OpenStudy (anonymous):

solve C=3(2-5h) for f in terms of c and h. show how you got the answer.

OpenStudy (anonymous):

c=3(2f-5h)

OpenStudy (ash2326):

C=3(2f-5h) multiply the term in bracket by 3 C=6f-15h tainh 15h to left side C+15h=6f dividing by 6 whole equation f=(c+15h)/6

OpenStudy (ash2326):

sorry it's take in 4th line

OpenStudy (anonymous):

Thats ok